# Towards Mobile Sensing with Event Cameras on High-mobility Resource-constrained Devices: A Survey

### v1

@@ Abstract
With the increasing complexity of mobile device applications, these devices are evolving toward high mobility. This shift imposes new demands on mobile sensing, particularly in terms of achieving high accuracy and low latency. Event-based vision has emerged as a disruptive paradigm, offering high temporal resolution, low latency, and energy efficiency, making it well-suited for high-accuracy and low-latency sensing tasks on high-mobility platforms. However, the presence of substantial noisy events, the lack of inherent semantic information, and the large data volume pose significant challenges for event-based data processing on resource-constrained mobile devices. This paper surveys the literature over the period 2014-2024, provides a comprehensive overview of event-based mobile sensing systems, covering fundamental principles, event abstraction methods, algorithmic advancements, hardware and software acceleration strategies. We also discuss key applications of event cameras in mobile sensing, including visual odometry, object tracking, optical flow estimation, and 3D reconstruction, while highlighting the challenges associated with event data processing, sensor fusion, and real-time deployment. Furthermore, we outline future research directions, such as improving event camera hardware with advanced optics, leveraging neuromorphic computing for efficient processing, and integrating bio-inspired algorithms to enhance perception. To support ongoing research, we provide an open-source Online Sheet with curated resources and recent developments. We hope this survey serves as a valuable reference, facilitating the adoption of event-based vision across diverse applications.
@@ I Introduction
Figure 1: Mobile devices are employed in a wide range of applications. Among these, the most critical steps involve estimating the devices’ states, perceiving surrounding environmental structures, and understanding the relationships between devices and their environment. As mobile devices advance toward high-mobility design, these mobile platforms impose new demands on mobile sensing, particularly in terms of high accuracy and low latency. This necessitates close coordination between sensors and sensing algorithms: $(i)$ Sensors must acquire high-accuracy data with low latency; $(ii)$ Sensing algorithms must process the data efficiently and accurately under constrained resources. Existing sensors fail to fully meet these requirements; Event cameras, capable of asynchronously capturing pixel-level intensity changes with $\mu s$-level latency, hold the potential to revolutionize mobile sensing. This survey provides a comprehensive review of event cameras and the development of efficient and accurate algorithms tailored to them.
@@
High mobility trend of mobile devices. In context of smart city scenarios, the demand for mobile devices is progressively increasing, with their mission profiles evolving toward execution of 4D tasks, which are characterized as deep, dull, dangerous, and dirty [20, 21, 22]. The expanding scale of cities also necessitates that mobile devices complete various tasks within shorter time frames, driving their evolution toward high-speed operation [1, 23, 24]. Consequently, the development of mobile devices is exhibiting new trends toward high mobility. For instance, DJI’s industrial inspection drones can cruise at speeds of 21 $m/s$ [25], while Wing’s delivery drones fly at 30 $m/s$ to deliver packages [26].
@@
New challenges for mobile sensing. As mobile devices evolve toward high mobility design, mobile sensing is required to advance toward high accuracy and low latency, enabling these mobile devices to perceive their state and surroundings in $millimeter$-level accuracy with $millisecond$-level latency, thereby facilitating faster responses and more precise adjustments. This evolution establishes new objectives for the sensors and data processing algorithms involved in mobile sensing tasks: $(i)$ On the sensor input front, it is essential to acquire higher accuracy raw data with lower latency. $(ii)$ On data processing algorithms front, efficient processing of sensor measurements is essential to enhance accuracy in mobile sensing tasks while optimizing performance on resource-constrained platforms.
@@
Existing sensors for mobile sensing. However, existing sensors are increasingly inadequate in meeting the high accuracy and low latency requirements for mobile sensing, particularly in environmental sensing and interaction for high-mobility devices. $(i)$ Radar-based solutions leverage sensors such as LiDAR [1, 27], mmWave radar [28, 29], or ultrasound radar [30], which emit signals within specific spectrums and calculate distances from reflected signals. These methods track changes in distances to update device locations and infer spatial relationships with surrounding objects. However, the high sensing delay, low sensing accuracy of these sensors render them poorly suited for high-mobility mobile devices. LiDAR can achieve millimeter-level accuracy, but it requires accumulating points into frames with a low output frequency (e.g., 10 $Hz$), causing a delay equal to the frame interval (typically 100 $ms$). Meanwhile, mmWave radar provides millisecond-level latency but suffers from limited spatial resolution, preventing millimeter-level sensing accuracy. $(ii)$ Camera-based methods utilize monocular [31, 32, 33] and stereo camera [34, 35] for self-state estimation and environmental understanding via SLAM techniques [36]. However, they are computationally intensive and constrained by limited temporal resolution (e.g., <30 $Hz$), high sensing latency (e.g., >30 $ms$), and standard dynamic range (e.g., 60 $dB$), making them unsuitable for high-mobility mobile platforms.
@@
New sensor: Event camera. The event cameras are novel bio-inspired sensors that outputs pixel-wise intensity changes in an asynchronous manner [42, 37]. Unlike frame cameras, it generates output based on scene dynamics rather than a global clock that is independent of the scene [43, 44]. The event cameras offer four key advantages that align well with the requirements of mobile sensing tasks for high-mobility mobile devices: $(i)$ The $\mu s$-level temporal resolution enables the capture of high-speed motions without motion blur, supporting accurate sensing during high-speed operations [45, 46]. $(ii)$ The $\mu s$-level sensing latency allows the report of environmental changes to mobile devices almost instantaneously [47]. $(iii)$ The high dynamic range (HDR), which is 140 $dB$ compared to 60 $dB$ for standard cameras, making it effective in diverse lighting conditions [48, 43]. $(iv)$ The low power consumption (e.g., 0.5 $W$) makes it particularly suitable for efficient designed mobile devices [49]. These advantages position event cameras as a promising technology to empower mobile devices designed for high-speed operation and efficiency.
@@
Event-based data processing algorithms. Although event cameras enable high-accuracy data acquisition with low-latency, their unique characteristics introduce three challenges to data processing: $(i)$ Sensitivity to illumination results in significant noise in the sensor output. $(ii)$ The lack of inherent semantic information leads to feature extraction failure. $(iii)$ The data volume is substantial (e.g., thousands of events may be generated within a short time), leading to excessive computation overhead for mobile devices. These challenges make efficient and accurate event-based data processing essential for resource-constrained, high-mobility mobile devices, allowing them to accurately perceive their state and understand their surroundings with minimal latency. This survey categorizes event-based data processing algorithms into six key phases and provides a comprehensive review of each: event representation, event-based denoising, event-based filtering and feature extraction, event-based matching, event-based mapping, and event-based hardware/software acceleration. For specific applications, several end-to-end learning algorithms have been proposed to fully exploit data-driven approaches for task optimization.
@@
Difference between existing surveys. As shown in Fig.2, unlike existing surveys that summarize the applications of event cameras across different domains [37, 38, 39, 40, 41], this survey focuses on how event cameras aid high-mobility resource-constrained mobile devices in high-accuracy and low-latency self-state estimation and environmental understanding, surveying the literature over the period 2014-2024. It outlines processing workflow of event data and reviews the latest advancements at each stage of this workflow. Using the key metrics of accuracy and efficiency in mobile computing, we summarize various methods in each stage to provide a deeper analysis of cutting-edge research. It also covers works on event-based hardware and software acceleration, offering critical insights for deploying event cameras on resource-constrained mobile devices. Finally, the survey discusses application scenarios of event cameras on mobile platforms.
@@
$\bullet$ We present a comprehensive review of how event cameras enhance high-mobility resource-constrained mobile devices by enabling high-accuracy, low-latency self-state estimation and environmental perception.
@@
Online resource. This survey presents a comprehensive review of event-based sensing systems, focusing on key technological advancements and practical applications. To further support the research community, we have established a open-source Online Sheet  [^1] Event-based mobile sensing resource that will be regularly updated, ensuring access to the latest developments and fostering continued innovation in event-based sensing system.
@@ II-C Advantages of event cameras in mobile sensing
High temporal resolution. The high mobility of mobile devices leads to rapid environmental changes, causing motion blur in standard frame cameras and making timely sensing about changes in environment and self-state difficult for radar-based or camera-based solution due to their low spatial-temporal resolution. Event cameras, with microsecond resolution and the ability to detect changes without motion blur, offer crucial support for quickly sensing changes in the environment and device state.
Low latency. The high mobility of mobile devices requires rapid awareness of environmental and self-state changes to respond effectively. Standard frame cameras depend on a global exposure time (about 20 $ms$) and additional processing time (10-20 $ms$), which delays reaction times and increases the risk of collisions. In contrast, event cameras operate with independent pixels that generate events immediately upon brightness changes, without requiring a global exposure time. This results in sub-millisecond latency, allowing mobile devices to detect changes and respond quickly.
@@ II-D Challenges of event cameras in mobile sensing
(i) How to mitigate event bursts and accurately extract features from event data, given high-speed operation of mobile devices and its lack of inherent semantic information? Event cameras are highly sensitive to illumination changes, with even minor variations triggering numerous events. On high-mobility mobile devices, rapid scene changes captured by onboard event cameras can trigger event bursts, generating thousands of events in a short time. Event bursts and the lack of inherent semantic information hinder the extraction of critical environmental information, as they are often overwhelmed by irrelevant events caused by device motion.
@@ IV-A Event-based denoising
Deep learning-based methods. Recent advances in deep learning have revolutionized event denoising, enabling automated and adaptive solutions. Deep learning models train on noisy and clean event pairs to learn mappings, reducing reliance on manual parameter tuning. Baldwin et al. [106] proposed the Event Denoising Convolutional Neural Network (EDnCNN), which integrates frames and IMU data to classify events as signal or noise. EventZoom [116] utilizes a U-Net architecture for efficient noise-to-noise denoising, achieving superior performance in handling noisy event streams. AEDNet [107] directly processes raw DVS data, preserving inherent spatio-temporal correlations while denoising. Additional methods, such as K-SVD [117] and multilayer perceptron-based denoising filters (MLPF) [104], focus on sparse feature extraction and event probability estimation, further advancing the capabilities of learning-based denoising frameworks.

### v2

# Towards Mobile Sensing with Event Cameras on High-agility Resource-constrained Devices: A Survey

@@ Abstract
With the increasing complexity of mobile device applications, these devices are evolving toward high agility. This shift imposes new demands on mobile sensing, particularly in terms of achieving high accuracy and low latency. Event-based vision has emerged as a disruptive paradigm, offering high temporal resolution, low latency, and energy efficiency, making it well-suited for high-accuracy and low-latency sensing tasks on high-agility platforms. However, the presence of substantial noisy events, the lack of inherent semantic information, and the large data volume pose significant challenges for event-based data processing on resource-constrained mobile devices. This paper surveys the literature over the period 2014-2024, provides a comprehensive overview of event-based mobile sensing systems, covering fundamental principles, event abstraction methods, algorithmic advancements, hardware and software acceleration strategies. We also discuss key applications of event cameras in mobile sensing, including visual odometry, object tracking, optical flow estimation, and 3D reconstruction, while highlighting the challenges associated with event data processing, sensor fusion, and real-time deployment. Furthermore, we outline future research directions, such as improving event camera hardware with advanced optics, leveraging neuromorphic computing for efficient processing, and integrating bio-inspired algorithms to enhance perception. To support ongoing research, we provide an open-source Online Sheet with curated resources and recent developments. We hope this survey serves as a valuable reference, facilitating the adoption of event-based vision across diverse applications.
@@ I Introduction
Figure 1: Mobile devices are employed in a wide range of applications. Among these, the most critical steps involve estimating the devices’ states, perceiving surrounding environmental structures, and understanding the relationships between devices and their environment. As mobile devices advance toward high-agility design, these mobile platforms impose new demands on mobile sensing, particularly in terms of high accuracy and low latency. This necessitates close coordination between sensors and sensing algorithms: $(i)$ Sensors must acquire high-accuracy data with low latency; $(ii)$ Sensing algorithms must process the data efficiently and accurately under constrained resources. Existing sensors fail to fully meet these requirements; Event cameras, capable of asynchronously capturing pixel-level intensity changes with $\mu s$-level latency, hold the potential to revolutionize mobile sensing. This survey provides a comprehensive review of event cameras and the development of efficient and accurate algorithms tailored to them.
@@
High agility trend of mobile devices. In context of smart city scenarios, the demand for mobile devices is progressively increasing, with their mission profiles evolving toward execution of 4D tasks, which are characterized as deep, dull, dangerous, and dirty [20, 21, 22]. The expanding scale of cities also necessitates that mobile devices complete various tasks within shorter time frames, driving their evolution toward high-speed operation [1, 23, 24]. Consequently, the development of mobile devices is exhibiting new trends toward high agility. For instance, DJI’s industrial inspection drones can cruise at speeds of 21 $m/s$ [25], while Wing’s delivery drones fly at 30 $m/s$ to deliver packages [26].
@@
New challenges for mobile sensing. As mobile devices evolve toward high agility design, mobile sensing is required to advance toward high accuracy and low latency, enabling these mobile devices to perceive their state and surroundings in $millimeter$-level accuracy with $millisecond$-level latency, thereby facilitating faster responses and more precise adjustments. This evolution establishes new objectives for the sensors and data processing algorithms involved in mobile sensing tasks: $(i)$ On the sensor input front, it is essential to acquire higher accuracy raw data with lower latency. $(ii)$ On data processing algorithms front, efficient processing of sensor measurements is essential to enhance accuracy in mobile sensing tasks while optimizing performance on resource-constrained platforms.
@@
Existing sensors for mobile sensing. However, existing sensors are increasingly inadequate in meeting the high accuracy and low latency requirements for mobile sensing, particularly in environmental sensing and interaction for high-agility devices. $(i)$ Radar-based solutions leverage sensors such as LiDAR [1, 27], mmWave radar [28, 29], or ultrasound radar [30], which emit signals within specific spectrums and calculate distances from reflected signals. These methods track changes in distances to update device locations and infer spatial relationships with surrounding objects. However, the high sensing delay, low sensing accuracy of these sensors render them poorly suited for high-agility mobile devices. LiDAR can achieve millimeter-level accuracy, but it requires accumulating points into frames with a low output frequency (e.g., 10 $Hz$), causing a delay equal to the frame interval (typically 100 $ms$). Meanwhile, mmWave radar provides millisecond-level latency but suffers from limited spatial resolution, preventing millimeter-level sensing accuracy. $(ii)$ Camera-based methods utilize monocular [31, 32, 33] and stereo camera [34, 35] for self-state estimation and environmental understanding via SLAM techniques [36]. However, they are computationally intensive and constrained by limited temporal resolution (e.g., <30 $Hz$), high sensing latency (e.g., >30 $ms$), and standard dynamic range (e.g., 60 $dB$), making them unsuitable for high-agility mobile platforms.
@@
New sensor: Event camera. The event cameras are novel bio-inspired sensors that outputs pixel-wise intensity changes in an asynchronous manner [42, 37]. Unlike frame cameras, it generates output based on scene dynamics rather than a global clock that is independent of the scene [43, 44]. The event cameras offer four key advantages that align well with the requirements of mobile sensing tasks for high-agility mobile devices: $(i)$ The $\mu s$-level temporal resolution enables the capture of high-speed motions without motion blur, supporting accurate sensing during high-speed operations [45, 46]. $(ii)$ The $\mu s$-level sensing latency allows the report of environmental changes to mobile devices almost instantaneously [47]. $(iii)$ The high dynamic range (HDR), which is 140 $dB$ compared to 60 $dB$ for standard cameras, making it effective in diverse lighting conditions [48, 43]. $(iv)$ The low power consumption (e.g., 0.5 $W$) makes it particularly suitable for efficient designed mobile devices [49]. These advantages position event cameras as a promising technology to empower mobile devices designed for high-speed operation and efficiency.
@@
Event-based data processing algorithms. Although event cameras enable high-accuracy data acquisition with low-latency, their unique characteristics introduce three challenges to data processing: $(i)$ Sensitivity to illumination results in significant noise in the sensor output. $(ii)$ The lack of inherent semantic information leads to feature extraction failure. $(iii)$ The data volume is substantial (e.g., thousands of events may be generated within a short time), leading to excessive computation overhead for mobile devices. These challenges make efficient and accurate event-based data processing essential for resource-constrained, high-agility mobile devices, allowing them to accurately perceive their state and understand their surroundings with minimal latency. This survey categorizes event-based data processing algorithms into six key phases and provides a comprehensive review of each: event representation, event-based denoising, event-based filtering and feature extraction, event-based matching, event-based mapping, and event-based hardware/software acceleration. For specific applications, several end-to-end learning algorithms have been proposed to fully exploit data-driven approaches for task optimization.
@@
Difference between existing surveys. As shown in Fig.2, unlike existing surveys that summarize the applications of event cameras across different domains [37, 38, 39, 40, 41], this survey focuses on how event cameras aid high-agility resource-constrained mobile devices in high-accuracy and low-latency self-state estimation and environmental understanding, surveying the literature over the period 2014-2024. It outlines processing workflow of event data and reviews the latest advancements at each stage of this workflow. Using the key metrics of accuracy and efficiency in mobile computing, we summarize various methods in each stage to provide a deeper analysis of cutting-edge research. It also covers works on event-based hardware and software acceleration, offering critical insights for deploying event cameras on resource-constrained mobile devices. Finally, the survey discusses application scenarios of event cameras on mobile platforms.
@@
$\bullet$ We present a comprehensive review of how event cameras enhance high-agility resource-constrained mobile devices by enabling high-accuracy, low-latency self-state estimation and environmental perception.
@@
Online resource. This survey presents a comprehensive review of event-based sensing systems, focusing on key technological advancements and practical applications. To further support the research community, we have established a open-source Online Sheet  [^1] Event-based mobile sensing resource, which is adapted from [50]. This online sheet will be regularly updated, ensuring access to the latest developments and fostering continued innovation in event-based sensing system.
@@ II-C Advantages of event cameras in mobile sensing
High temporal resolution. The high agility of mobile devices leads to rapid environmental changes, causing motion blur in standard frame cameras and making timely sensing about changes in environment and self-state difficult for radar-based or camera-based solution due to their low spatial-temporal resolution. Event cameras, with microsecond resolution and the ability to detect changes without motion blur, offer crucial support for quickly sensing changes in the environment and device state.
Low latency. The high agility of mobile devices requires rapid awareness of environmental and self-state changes to respond effectively. Standard frame cameras depend on a global exposure time (about 20 $ms$) and additional processing time (10-20 $ms$), which delays reaction times and increases the risk of collisions. In contrast, event cameras operate with independent pixels that generate events immediately upon brightness changes, without requiring a global exposure time. This results in sub-millisecond latency, allowing mobile devices to detect changes and respond quickly.
@@ II-D Challenges of event cameras in mobile sensing
(i) How to mitigate event bursts and accurately extract features from event data, given high-speed operation of mobile devices and its lack of inherent semantic information? Event cameras are highly sensitive to illumination changes, with even minor variations triggering numerous events. On high-agility mobile devices, rapid scene changes captured by onboard event cameras can trigger event bursts, generating thousands of events in a short time. Event bursts and the lack of inherent semantic information hinder the extraction of critical environmental information, as they are often overwhelmed by irrelevant events caused by device motion.
@@ IV-A Event-based denoising
Deep learning-based methods. Recent advances in deep learning have revolutionized event denoising, enabling automated and adaptive solutions. Deep learning models train on noisy and clean event pairs to learn mappings, reducing reliance on manual parameter tuning. Baldwin et al. [107] proposed the Event Denoising Convolutional Neural Network (EDnCNN), which integrates frames and IMU data to classify events as signal or noise. EventZoom [117] utilizes a U-Net architecture for efficient noise-to-noise denoising, achieving superior performance in handling noisy event streams. AEDNet [108] directly processes raw DVS data, preserving inherent spatio-temporal correlations while denoising. EDformer [118] utilizes an event-by-event denoising model based on a transformer model. Additional methods, such as K-SVD [119] and multilayer perceptron-based denoising filters (MLPF) [105], focus on sparse feature extraction and event probability estimation, further advancing the capabilities of learning-based denoising frameworks.